OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 3x

OpenStudy (anonymous):

Cos3x = Cos (2x + x) = Cos2xCosx - Sin2xSinx

OpenStudy (anonymous):

where did you get the -sin2xsixn?

OpenStudy (anonymous):

This is from cos(a + b) = cosAcosB - sinAsinB. This is an identity.

OpenStudy (anonymous):

okay

OpenStudy (kagıtucak):

cos(2x+x)=cos(2x).cos(x)-sin(2x).sinx you can put 2cos^2(x)-1 or 1-2sin^2(x) instead of cos(2x) then ı chose the first one doesnt matter. =2cos^3(x)-cos(x)-2.sin(x).cos(x).cos(x)=cos(x)[2cos^2(x)-1-2sin(x)cos(x)]

OpenStudy (anonymous):

These are the answer choices cos x - 4 cos x sin2x -sin3x + 2 sin x cos x -sin2x + 2 sin x cos x 2 sin2x cos x - 2 sin x cos x